OpenStudy (anonymous):

how do i solve dis if 5/2x+2=10 then x=

OpenStudy (ivettef365):

5/2x = 10 - 2 5/2x = 8 then multiply by 2x and you have 5 = 16x then divide by 16 and you have 5/16 = x

OpenStudy (anonymous):

Isolate the x. First you take 2 away from both sides. This leaves you with 5/2x=8. Then multiply both sides by 2x, so now we have 5=16x. Finally we solve for x by dividing both sides by 16 in order to isolate the x. Our answer would be 5/16=x

OpenStudy (anonymous):

Oops sorry c: Just saw someone else answered

OpenStudy (ivettef365):

Is it like this 5 -- + 2 = 10 2x

OpenStudy (whpalmer4):

Is that \[\frac{5}{2x+2} = 10\]or\[\frac{5}{2x}+2=10\]?

OpenStudy (whpalmer4):

If the former, multiply both sides by (2x+2). Then distribute the right hand side and solve for x.

OpenStudy (anonymous):

NO the x is seperate (ivettef365)

OpenStudy (anonymous):

Then look at the first response by @ivettef365 he answered it correctly

OpenStudy (whpalmer4):

We have to ask because all too many people will write 5/2x+2 = 10 when they mean 5/(2x+2) = 10

OpenStudy (anonymous):

5 -- + 2 = 10 this is correct help me solve please 2x

OpenStudy (anonymous):

Isolate the x. First you take 2 away from both sides. This leaves you with 5/2x=8. Then multiply both sides by 2x, so now we have 5=16x. Finally we solve for x by dividing both sides by 16 in order to isolate the x. Our answer would be 5/16=x

OpenStudy (whpalmer4):

@ivettef365 did that one correctly: \[\frac{5}{2x}+2 = 10\]Subtract 2 from both sides:\[\frac{5}{2x}+2-2 = 10-2\]\[\frac{5}{2x}=8\]Multiply both sides by \(2x\)\[\frac{5}{2x}*2x=8*2x\]\[5=16x\]Now divide both sides by 16 to solve for \(x\):\[\frac{5}{16} = \frac{\cancel{16}x}{\cancel{16}} = x\]